OpenStudy (anonymous):

Tony has two concentrations of antifreeze in his garage 50% and 100%. In most climates, a 50% concentration is optimal, but Tony lives in Minnesota and wants a 65% concentration of antifreeze in his car. The cooling system holds 13 quarts. How many quarts of each type of antifreeze should he use to fill his system with a 65% concentration?

OpenStudy (radar):

Somehow I am missing the boat. I keep coming up with an 85% solution rather than 65%. Please review my work and find the error. Let X = number of quarts of the 50% concentrate needed. Let Y = number of quarts of the 100% solution (pure) X + Y = 13 Any thing wrong so far?????

OpenStudy (radar):

The amount of antifreeze for a 65% solution in a 13 quart radiator would be: (.65) x 13 = 8.45 quarts or 8.45 quarts of pure antifreeze (100%) Does this make sense? By the way I feel sorry for Tony in Minnesota.

OpenStudy (radar):

Using the variable assignments (labels) in the first post then we can state this equation. .5X + Y =8.45 qts (pure antifreeze).

OpenStudy (radar):

We now have two equations, two unknowns thus it can be solved. Taking the equation X + Y = 13 State in terms of X getting X= 13-Y substitute this in our other equation. .5(13-Y) + Y = 8.45 this becomes 6.5 - .5Y + Y = 8.45 .5Y = 8.45 - 6.5 .5Y =1.95 Y = 3.9 qts of the 100% solution substuting in the X + Y = 13 we solve for X X=13-3.9 =9.1 quarts of the 50% solution. So far looking correct, but we now need to verify.

OpenStudy (radar):

X+Y = 13............9.1 + 3.9 = 13 yup (.5)(9.1) + 3.9 =8.45.......4.55+3.9 =8.45 yup but is the solution 65% what do you think?

OpenStudy (radar):

It checks out o.k. Those are the correct answers.

OpenStudy (anonymous):

Thanks